# The Diophantine problem in isotropic reductive groups

### v1

@@ 3 Technical lemmas
A pair of subsets $(X,X_{0})\subseteq(M,M_{0})$ generates $(M,M_{0})$ if $X_{0}$ generates the $K$-module $M_{0}$ and $X$ generates the $K$-module $M/M_{0}$.
@@
Every locally free $2$-step nilpotent $K$-module $(M,M_{0})$ determines a representable fpqc sheaf $\mathbb{M}_{0}(E)=E\otimes_{K}M_{0}$, $\mathbb{M}(E)=M\boxtimes E=E\otimes_{K}M_{0}\mathbin{\dot{\oplus}}E\otimes_{K}M_{1}$ of locally free $2$-step nilpotent modules (where $K\to E$ is a ring homomorphism), the latter one is independent on the choice of the splitting [8, §4]. Also, locally free $2$-step nilpotent $K$-modules satisfy the fpqc descent. If $G$ is a group scheme acting on such a sheaf $(\mathbb{M},\mathbb{M}_{0})$ by automorphisms of $2$-step nilpotent modules and $G$ stabilizes some generating set $(X,X_{0})\subseteq(M,M_{0})$, then the action is trivial.
By [9, §2] there are canonical homomorphisms $t_{\alpha}\colon\mathfrak{g}_{\alpha}\to G_{\alpha}(K)$ for non-ultrashort $\alpha\in\Phi$ and $t_{\alpha}\colon\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to G_{\alpha}(K)$ for ultrashort $\alpha$. In the second case $\mathfrak{g}_{\alpha}\dotplus\mathfrak{g}_{2\alpha}$ is a locally free $2$-step nilpotent $K$-module. Since $G$ is root graded, the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument if $\alpha$ is long and $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$, i.e. $\mathfrak{g}_{\beta}\to\mathrm{Hom}(\mathfrak{g}_{\alpha},\mathfrak{g}_{\alpha+\beta})$ is injective. There are also several other instances of non-degeneracy, we check them case by case.
@@
In the orthogonal-like case $\mathfrak{g}_{\alpha}\cong R$ for a ring $R$ (an Azumaya algebra over $K$ or its quadratic étale extension) and $\mathfrak{g}_{\beta}\cong M_{R}$ for some right faithfully projective $A$-module, the pairing is just $M\times R\to M,\,(m,r)\mapsto mr$ up to an isomorphism. In the symplectic-like case $\mathfrak{g}_{\beta}\cong R$ for a ring $R$ with a $\lambda$-involution and $\mathfrak{g}_{\alpha}\cong\Lambda$ for a form parameter $\Lambda\leq R$, the pairing is $R\times\Lambda\to R,\,(r,u)\mapsto ru$ up to an isomorphism. This pairing is also non-degenerate, the intersection $\Lambda\cap R^{*}$ is non-empty because $G$ is root graded. ∎
@@
If $G$ is classical, then this follows from the non-degeneracy of the hermitian form on the defining representation. Otherwise there are the cases ${{}^{2}\!{\mathrm{E}_{6,2}^{16^{\prime}}}}$, $\mathrm{E}_{7,2}^{31}$, and $\mathsf{E}_{8,2}^{66}$. It is possible to check using the root diagrams of $\mathsf{E}_{\ell}$ that the required Lie bracket is non-degenerate. ∎
@@ 4 Some normalizers and centralizers
Let $G$ be an isotropic reductive group over $K$ with root system $\Phi$ of rank $\geq 2$. Let also $\alpha\in\Phi$ be a long root and $X\subseteq\mathfrak{g}_{\alpha}$ be a generating set of a $K$-module. Then the group scheme
@@
Clearly, $P\leq N$ and we may assume that $K$ is local. By lemma 1 it suffices to check that $G_{\Sigma}\cap N=1$, where $\Sigma=\{\beta\mid\angle(\alpha,\beta)>\frac{\pi}{2}\}$. Choose a linear map $f\colon\mathbb{R}\Phi\to\mathbb{R}$ such that $\operatorname{Ker}(f)\cap\Phi=\mathbb{R}\alpha\cap\Phi$ and let $\Phi_{s}=\{\beta\in\Phi\mid\mathrm{sign}(f(\beta))=s\}$, $\Sigma_{s}=\Phi_{s}\cap\Sigma$ for $s\in\{{-},0,{+}\}$, so $\Phi_{\pm}$ are unipotent subsets.
Take any element $g=g_{-}g_{0}g_{+}\in G_{\Sigma}\cap N$ with $g_{s}\in G_{\Sigma_{s}}$, then $g_{\pm}=\prod_{\beta\in\Sigma_{\pm}\setminus 2\Sigma_{\pm}}t_{\beta}(y_{\beta})$ for any linear orders on $\Sigma_{\pm}\setminus 2\Sigma_{\pm}$ and $g_{0}\in G_{-\alpha}$ (or $g_{0}\in G_{-\alpha/2}$ if $\Phi$ is of type $\mathsf{BC}_{\ell}$). For any $x\in X$ there is $x^{\prime}\in\mathfrak{g}_{\alpha}$ such that $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. Comparing both sides of $g_{-}\,g_{0}\,t_{\alpha}(x)\,g_{+}^{t_{\alpha}(x)}={{}^{t_{\alpha}(x^{\prime})}\!{g_{-}}}\,t_{\alpha}(x^{\prime})\,g_{0}\,g_{+}$ we get $[g_{+},t_{\alpha}(x)]=1$ and $[g_{-},t_{\alpha}(x^{\prime})]=1$. Since $x$ and $x^{\prime}$ runs generating sets of $\mathfrak{g}_{\alpha}$ and $G$ is root graded (so the commutator map $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument for $\beta\in\Sigma_{\pm}$) both $g_{+}$ and $g_{-}$ are trivial.
It follows that $g=g_{0}\in G_{-\alpha^{\prime}}$ for $\alpha^{\prime}\in\{\alpha,\frac{\alpha}{2}\}$. Now suppose that $\Phi$ is not of the type $\mathsf{BC}_{\ell}$ and take a root $\beta\in\Phi$ such that $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$ and $\beta$ is long if $\Phi$ is of the type $\mathsf{G}_{2}$. Conjugate $t_{\beta}(z)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. We get $\bigl{[}\log g,[x,z]\bigr{]}=1$, where $[{-},{=}]\colon\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is the commutator map and $\log g\in\mathfrak{g}_{-\alpha}$ is the element such that $g=t_{-\alpha}(\log g)$. Since $G$ is root graded, $\mathfrak{g}_{\alpha+\beta}=[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]$, so $[\log g,\mathfrak{g}_{\alpha+\beta}]=0$ and $\log g=0$ (by lemma 3 if $\beta$ is short), i.e. $g=1$.
Finally, suppose that $\Phi$ is of type $\mathsf{BC}_{\ell}$. Take an ultrashort root $\beta$ orthogonal to $\alpha$ and conjugate $t_{\beta}(u)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. We get $[t_{\alpha}(x^{\prime}),[g,t_{\beta}(u)]]=1$, so $[g,t_{\beta}(u)]=1$ and $g\in G_{-\alpha}$ by lemma 4. In other words, we reduce to the previous case of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. ∎
@@
is a sheaf of locally free $2$-step nilpotent modules. The action of the group scheme $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$ on it commutes with the operations of $2$-step nilpotent modules, this may be easily checked by passing to a split form of $G$.
@@
$C=\{g\in G^{0}_{\Phi_{0}}\mid[g,t_{\alpha}(x_{\alpha,i})]=1\text{ for }f(\alpha)>0\}$
coincides with the scheme center $\mathrm{C}(G)\leq\mathrm{C}(L)$, where $\Phi_{0}=\Phi\cap\operatorname{Ker}(f)$ and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}$ are $K$-module generators if $\alpha$ is not ultrashort and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ generate the $K$-module $\mathfrak{g}_{\alpha}$ otherwise.
@@
Clearly, $C$ is the scheme centralizer of $G_{\Phi_{>0}}$ in $G^{0}_{\Phi_{0}}$, where $\Phi_{>0}=\{\alpha\in\Phi\mid f(\alpha)>0\}$. We may assume that $G$ splits and $\Phi=\widetilde{\Phi}$. Firstly, let us show that $C$ is contained in the maximal torus $T=L$. By theorem 1 and lemma 2 the group scheme $C$ is contained in $G^{0}_{\Sigma}$, where $\Sigma$ consists of the roots $\alpha\in\Phi_{0}$ such that $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all long roots $\beta\in\Phi_{>0}$. By considering rank $2$ root subsystems it may be shown that roots of $\Sigma$ are actually orthogonal to long roots from $\Phi_{>0}$.
Now if the set of long roots of $\Phi$ is indecomposable root system (of full rank), then the set of long roots from $\Phi\setminus\Phi_{0}$ generates this root subsystem e.g. by [9, lemma 1]. In this case necessarily $\Sigma=\varnothing$. Otherwise $\Phi=\mathsf{C}_{\ell}$, possibly $\mathsf{B}_{2}=\mathsf{C}_{2}$. Up to the choice of the base $\Delta\subseteq\Phi$ any $5$-grading of $\mathsf{C}_{\ell}$ maps basic roots to 00 except one, and this distinguished root is mapped to $1$ (if it is short) or $2$ (if it is long). It easily follows that $\Sigma=\mathsf{C}_{m}\subseteq\mathsf{C}_{\ell}$ is a proper root subsystem generated by a connected Dynkin subdiagram containing the long basic root or it is empty. But the representation of $\mathrm{Sp}(2m,K)$ on $G_{\{\alpha\in\mathsf{C}_{m+1}\mid\alpha\cdot\mu>0\}}/G_{\mu}$ is the defining representation of symplectic group scheme, where $\mu\in\mathsf{C}_{m+1}$ is the highest root. It follows that $C\leq T$ in all cases.
@@
If $\Phi$ is of type $\mathsf{BC}_{\ell}$ let $\mathrm{C}^{\mathrm{us}}(G)=\bigcap_{\alpha\in\mathsf{C}_{\ell}}\operatorname{C}_{L}(G_{\alpha})$ be the scheme centralizer of non-ultrashort root subgroups. This group scheme clearly contains the center. It turns out that it is reductive or finite of multiplicative type, the second case holds only for $\mathsf{E}_{8,2}^{66}$. The derived subgroup of $\mathrm{C}^{\mathrm{us}}(G)$ corresponds to the odd component of $u^{-1}(0)$ in the classical cases and to the component $\mathsf{A}_{1}$ in the case $\mathsf{E}_{7,2}^{31}$. More precisely, assuming that $G$ splits and $\operatorname{C}(G)=1$ we have
@@
${{}^{2}\!{\mathsf{A}_{n,r}^{(d)}}}$ $\mathbb{GL}_{n+1-2rd}$ $d\mid n+1$, $n\geq 1$, $2rd\leq n$, $r\geq 2$
$\mathsf{C}_{n,r}^{(d)}$ $\mathbb{S}\mathrm{p}_{2n-2rd}$ $2\leq d=2^{k}\mid 2n$, $rd<n$, $r\geq 2$
${{}^{1}\!{\mathsf{D}_{n,r}^{(d)}}}$, ${{}^{2}\!{\mathsf{D}_{n,r}^{(d)}}}$ $\mathbb{SO}_{2n-2rd}$ $r\geq 2$, $2\leq d=2^{k}\mid 2n$, $rd<n$
${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ $\mathbb{G}_{\mathrm{m}}$ none
$\mathsf{E}_{7,2}^{31}$ $\mathbb{SL}_{2}$ none
$\mathsf{E}_{8,2}^{66}$ $\mu_{2}$ none
In the classical cases the group scheme $\operatorname{C}^{\mathrm{us}}(G)$ can be computed using the block structure on $G$ as a matrix group. In the exceptional cases we apply lemma 1 to the splitting isotropic structure (the pinning) on $G$ and get $\operatorname{C}^{\mathrm{us}}(G)\leq T$ (for ${{}^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}}$ and $\mathsf{E}_{8,2}^{66}$) or $\mathbb{SL}_{2}\leq\operatorname{C}^{\mathrm{us}}(G)\leq T\,\mathbb{SL}_{2}$. The intersection $T\cap\operatorname{C}^{\mathrm{us}}(G)$ is easy to compute using the root diagrams.
For any root $\alpha\in\Phi$ let $\Gamma_{\alpha}=\{\beta\in\Phi\mid\alpha+\beta\notin\Phi\cup\{0\}\}$ and $Z_{\alpha}=\bigcup_{\beta\in\Gamma_{\alpha}}t_{\beta}(X_{\beta})\subseteq G(K)$ for some generating sets $X_{\beta}\subseteq\mathfrak{g}_{\beta}$ if $\beta$ is not ultrashort and generating sets $(X_{\beta},X_{2\beta})\subseteq(\mathfrak{g}_{\beta}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\beta},\mathfrak{g}_{2\beta})$ for ultrashort $\beta$. It is easy to see by considering all rank $2$ and rank $3$ irreducible root systems that $\Gamma_{\alpha}$ is closed.
@@
  • If the type of $\Phi$ is $\mathsf{BC}_{\ell}$ and $\alpha$ is ultrashort, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)\,G_{\alpha}$.
  • If $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ (using the convention $\mathrm{e}_{-i}=-\mathrm{e}_{i}$) is short and the type of $\Phi$ is $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, including $\mathsf{B}_{2}=\mathsf{C}_{2}$, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{2\mathrm{e}_{i}}\,G_{\mathrm{e}_{i}+\mathrm{e}_{j}}\,G_{2\mathrm{e}_{j}}$.
  • Otherwise $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,G_{\alpha}$.
@@
Considering the saturated root subsystem of rank $2$ (or rank $3$ in the doubly laced cases) containing $\alpha$ and arbitrary $\beta\in\Sigma\setminus\mathbb{R}\alpha$ we see that

### v2

@@ 3 Technical lemmas
A pair of subsets $(X,X_{0})\subseteq(M,M_{0})$ (i.e. $X\subseteq M$ and $X_{0}\subseteq M_{0}$) generates $(M,M_{0})$ if $X_{0}$ generates the $K$-module $M_{0}$ and $X$ generates the $K$-module $M/M_{0}$.
@@
Every locally free $2$-step nilpotent $K$-module $(M,M_{0})$ determines a representable fpqc sheaf $\mathbb{M}_{0}(E)=E\otimes_{K}M_{0}$, $\mathbb{M}(E)=M\boxtimes E=E\otimes_{K}M_{0}\mathbin{\dot{\oplus}}E\otimes_{K}M_{1}$ of locally free $2$-step nilpotent modules (where $K\to E$ is a ring homomorphism), the latter one is independent on the choice of the splitting [25, §4]. Also, locally free $2$-step nilpotent $K$-modules satisfy the fpqc descent. If $G$ is a group scheme acting on such a sheaf $(\mathbb{M},\mathbb{M}_{0})$ by automorphisms of $2$-step nilpotent modules and $G$ stabilizes some generating set $(X,X_{0})\subseteq(M,M_{0})$, then the action is trivial.
By [26, §2] there are canonical homomorphisms $t_{\alpha}\colon\mathfrak{g}_{\alpha}\to G_{\alpha}(K)$ for non-ultrashort $\alpha\in\Phi$ and $t_{\alpha}\colon\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to G_{\alpha}(K)$ for ultrashort $\alpha$. In the second case $\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ means a locally free $2$-step nilpotent $K$-module with the canonical short exact sequence $0\to\mathfrak{g}_{\alpha}\to\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}\to\mathfrak{g}_{2\alpha}\to 0$, though its splitting is non-canonical. Since $G$ is root graded locally in Zariski topology (by our additional assumption on isotropic data), the Lie bracket $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument if $\alpha$ is long and $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$, i.e. $\mathfrak{g}_{\beta}\to\mathrm{Hom}(\mathfrak{g}_{\alpha},\mathfrak{g}_{\alpha+\beta})$ is injective. There are also several other instances of non-degeneracy, we check them case by case.
@@
Without loss of generality, $G$ splits. For the Tits indices $\mathsf{B}_{n,2}$ and $^{k}\!{\mathsf{D}_{n,r}^{(1)}}$} we have $\mathfrak{g}_{\alpha}\cong K$ and $\mathfrak{g}_{\beta}\cong K^{n}$, the pairing is isomorphic to $K^{n}\times K\to K^{n},\,(m,x)\mapsto mx$.
For the Tits indices $\mathsf{C}_{2d,2}^{(d)}$ or $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} we have $\mathfrak{g}_{\beta}\cong\operatorname{M}(d,K)$ and
$\mathfrak{g}_{\alpha}\cong\Lambda=\{x\in\operatorname{M}(d,K)\mid x^{\mathrm{t}}=\pm x\},$
the pairing is isomorphic to $\operatorname{M}(d,K)\times\Lambda\to\operatorname{M}(d,K),\,(x,u)\mapsto xu$. This pairing is also non-degenerate because the sign - appears in $\Lambda$ for the Tits index $^{1}\!{\mathsf{D}_{2d,2}^{(d)}}$} and in this case $d\geq 2$.
Finally, if the Tits index is $^{2}\!{\mathsf{A}_{4d-1,2}^{(d)}}$}, then $\mathfrak{g}_{\beta}\cong\operatorname{M}(d,K)\times\operatorname{M}(d,K)$, $\mathfrak{g}_{\alpha}\cong\operatorname{M}(d,K)$, and the pairing is isomorphic to
$(\operatorname{M}(d,K)\times\operatorname{M}(d,K))\times\operatorname{M}(d,K)\to\operatorname{M}(d,K)\times\operatorname{M}(d,K),\,((x,y),u)\mapsto(xu,-uy).\qed$
@@
In the classical cases fppf locally the pairing is isomorphic to the product map
$\operatorname{M}(d,n,K)\times\operatorname{M}(n,d,K)\to\operatorname{M}(d,K)$
or to a pair of such maps. The exceptional cases $^{2}\!{\mathrm{E}_{6,2}^{16^{\prime}}}$}, $\mathrm{E}_{7,2}^{31}$, and $\mathsf{E}_{8,2}^{66}$ can be checked using [22]. ∎
@@ 4 Some normalizers and centralizers
Let $G$ be an isotropic reductive group over $K$ with root system $\Phi$ of rank $\geq 2$. Let also $\alpha\in\Phi$ be a long root and $X\subseteq\mathfrak{g}_{\alpha}$ be a generating set of a $K$-module. Then the group subscheme
@@
Clearly, $P\leq N$ and we may assume that $K$ is local. By lemma 1 it suffices to check that $G_{\Sigma}(K)\cap N(K)=1$, where $\Sigma=\{\beta\mid\angle(\alpha,\beta)>\frac{\pi}{2}\}$. Choose a linear map $f\colon\mathbb{R}\Phi\to\mathbb{R}$ such that $\operatorname{Ker}(f)\cap\Phi=\mathbb{R}\alpha\cap\Phi$ and let $\Phi_{s}=\{\beta\in\Phi\mid\mathrm{sign}(f(\beta))=s\}$, $\Sigma_{s}=\Phi_{s}\cap\Sigma$ for $s\in\{{-},0,{+}\}$, so $\Phi_{\pm}$ are unipotent subsets.
Take any element $g=g_{-}g_{0}g_{+}\in G_{\Sigma}(K)\cap N(K)$ with $g_{s}\in G_{\Sigma_{s}}$, then $g_{\pm}=\prod_{\beta\in\Sigma_{\pm}\setminus 2\Sigma_{\pm}}t_{\beta}(y_{\beta})$ for any linear orders on $\Sigma_{\pm}\setminus 2\Sigma_{\pm}$ and $g_{0}\in U_{-\alpha}(K)$ (or $g_{0}\in U_{-\alpha/2}(K)$ if $\Phi$ is of type $\mathsf{BC}_{\ell}$). For any $x\in X$ there is $x^{\prime}\in\mathfrak{g}_{\alpha}$ such that $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. Comparing both sides of $g_{-}\,g_{0}\,t_{\alpha}(x)\,g_{+}^{t_{\alpha}(x)}={{}^{t_{\alpha}(x^{\prime})}\!{g_{-}}}\,t_{\alpha}(x^{\prime})\,g_{0}\,g_{+}$ we get $g_{+},t_{\alpha}(x)]=1$ and $g_{-},t_{\alpha}(x^{\prime})]=1$. Since $x$ runs over a generating set of $\mathfrak{g}_{\alpha}$ and the commutator map $\mathfrak{g}_{\alpha}\times\mathfrak{g}_{\beta}\to\mathfrak{g}_{\alpha+\beta}$ is non-degenerate on the second argument for $\beta\in\Sigma_{\pm}$, $g_{+}$ is trivial, so $g_{-}$ is also trivial by changing the sign of $f$.
It follows that $g=g_{0}\in U_{-\alpha^{\prime}}(K)$ for $\alpha^{\prime}\in\{\alpha,\frac{\alpha}{2}\}$. Now suppose that $\Phi$ is not of the type $\mathsf{BC}_{\ell}$ and take a root $\beta\in\Phi$ such that $\frac{\pi}{2}<\angle(\alpha,\beta)<\pi$ and $\beta$ is long if $\Phi$ is of the type $\mathsf{G}_{2}$. Conjugate $t_{\beta}(z)$ by both sides of $g\,t_{\alpha}(x)=t_{\alpha}(x^{\prime})\,g$. We get $\bigl[\log g,[x,z]\bigr]=0$, where $\log g\in\mathfrak{g}_{-\alpha}$ is the element such that $g=t_{-\alpha}(\log g)$. Since $\mathfrak{g}_{\alpha+\beta}=[\mathfrak{g}_{\alpha},\mathfrak{g}_{\beta}]$, $\log g,\mathfrak{g}_{\alpha+\beta}]=0$ and $\log g=0$ (by lemma 3 if $\beta$ is short), i.e. $g=1$.
Finally, suppose that $\Phi$ is of type $\mathsf{BC}_{\ell}$. Take an ultrashort root $\beta$ orthogonal to $\alpha$ and conjugate $t_{\beta}(u)$ by both sides of $t_{\alpha}(x)\,g^{-1}=g^{-1}\,t_{\alpha}(x^{\prime})$. We get $t_{\alpha}(x^{\prime}),[g^{-1},t_{\beta}(u)]]=1$, so $g^{-1},t_{\beta}(u)]=1$ and $g\in U_{-\alpha}(K)$ by lemma 4. In other words, we reduce to the already considered case of $\mathsf{C}_{\ell}\subseteq\mathsf{BC}_{\ell}$. ∎
@@
is a sheaf of locally free $2$-step nilpotent modules. The action of the group scheme $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$ on it commutes with the operations of $2$-step nilpotent modules, this can be easily checked by passing to a split form of $G$.
@@
$C=\bigl\{g\in G^{0}_{\Phi\cap\operatorname{Ker}(f)}\mid[g,t_{\alpha}(x_{\alpha,i})]=1\text{ for all }\alpha\text{ with }f(\alpha)>0\bigr\}$
coincides with the scheme center $\mathrm{C}(G)\leq\mathrm{C}(L)$, $x_{\alpha,i}\in\mathfrak{g}_{\alpha}$ are $K$-module generators if $\alpha$ is not ultrashort and $x_{\alpha,i}\in\mathfrak{g}_{\alpha}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\alpha}$ generate the $K$-module $\mathfrak{g}_{\alpha}$ otherwise.
@@
Clearly, $C$ is the scheme centralizer of $G_{\Phi_{>0}}$ in $G^{0}_{\Phi\cap\operatorname{Ker}(f)}$, where $\Phi_{>0}=\{\alpha\in\Phi\mid f(\alpha)>0\}$. We can assume that $G$ splits and $\Phi=\widetilde{\Phi}$. Firstly, let us show that $C$ is contained in the maximal torus $T=L$. By theorem 1 and lemma 2 the group scheme $C$ is contained in $G^{0}_{\Sigma}$, where $\Sigma$ consists of the roots $\alpha\in\Phi_{0}$ such that $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all long roots $\beta\in\Phi_{>0}$. By considering rank $2$ root subsystems it follows that roots of $\Sigma$ are actually orthogonal to long roots from $\Phi_{>0}$.
Now if the set of long roots of $\Phi$ is an indecomposable root system of full rank, then the set of long roots from $\Phi\setminus\operatorname{Ker}(f)$ generates this root subsystem e.g. by [26, lemma 1]. In this case necessarily $\Sigma=\varnothing$. Otherwise $\Phi=\mathsf{C}_{\ell}$, possibly $\mathsf{B}_{2}=\mathsf{C}_{2}$. Up to the choice of the base $\Delta\subseteq\Phi$ any $5$-grading of $\mathsf{C}_{\ell}$ maps basic roots to 0 except one, and this distinguished root is mapped to $1$ (if it is short) or $2$ (if it is long). It easily follows that $\Sigma=\mathsf{C}_{m}\subseteq\mathsf{C}_{\ell}$ is a proper root subsystem generated by a connected Dynkin subdiagram containing the long basic root or it is empty. But the representation of $\mathrm{Sp}(2m,K)$ on $G_{\{\alpha\in\mathsf{C}_{m+1}\mid\alpha\cdot\mu>0\}}/U_{\mu}$ is the defining representation of symplectic group scheme, where $\mu\in\mathsf{C}_{m+1}$ is the highest root. It follows that $C\leq T$ in all cases.
@@
If $\Phi$ is of type $\mathsf{BC}_{\ell}$ with $\ell\geq 2$, let $\mathrm{C}^{\mathrm{us}}(G)=\bigcap_{\alpha\in\mathsf{C}_{\ell}}\operatorname{C}_{L}(U_{\alpha})$ be the scheme centralizer of non-ultrashort root subgroups. This group scheme clearly contains the center and it is contained in $L$ by theorem 1. It turns out that it is reductive or finite of multiplicative type, the second case holds only for $\mathsf{E}_{8,2}^{66}$. The derived subgroup of $\mathrm{C}^{\mathrm{us}}(G)$ corresponds to the “middle” component of $u^{-1}(0)$ in the classical cases and to the component $\mathsf{A}_{1}$ in the case $\mathsf{E}_{7,2}^{31}$. More precisely, assuming that $G$ is split and adjoint connected, we have the following.
@@
$^{2}\!{\mathsf{A}_{n,r}^{(d)}}$} $\mathbb{GL}_{n+1-2rd}$ $d\mid n+1$, $2rd\leq n$, $d\geq 1$, $r\geq 2$
$\mathsf{C}_{n,r}^{(d)}$ $\mathbb{S}\mathrm{p}_{2n-2rd}$ $2\leq d=2^{k}\mid 2n$, $rd\leq n-1$, $r\geq 2$
$^{1}\!{\mathsf{D}_{n,r}^{(d)}}$}, $^{2}\!{\mathsf{D}_{n,r}^{(d)}}$} $\mathbb{SO}_{2n-2rd}$ $r\geq 2$, $2\leq d=2^{k}\mid 2n$, $rd\leq n-1$
$^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} $\mathbb{G}_{\mathrm{m}}$
$\mathsf{E}_{7,2}^{31}$ $\mathbb{SL}_{2}$
$\mathsf{E}_{8,2}^{66}$ $\mu_{2}$
In the classical cases the group scheme $\operatorname{C}^{\mathrm{us}}(G)$ can be computed using the block structure on a finite central extension of $G$ as a matrix group. In the exceptional cases we apply theorem 1 to the splitting isotropic structure on $G$ and get $\operatorname{C}^{\mathrm{us}}(G)\leq T$ (for $^{2}\!{\mathsf{E}_{6,2}^{16^{\prime}}}$} and $\mathsf{E}_{8,2}^{66}$) or $\mathbb{SL}_{2}\leq\operatorname{C}^{\mathrm{us}}(G)\leq T\,\mathbb{SL}_{2}$, where $T$ is the maximal torus from the splitting isotropic structure. The intersection $T\cap\operatorname{C}^{\mathrm{us}}(G)$ is easy to compute using the root diagrams [22].
For any root $\alpha\in\Phi$ let $\Gamma_{\alpha}=\{\beta\in\Phi\mid\alpha+\beta\notin\Phi\cup\{0\}\}$ and $Z_{\alpha}=\bigcup_{\beta\in\Gamma_{\alpha}}t_{\beta}(X_{\beta})\subseteq G(K)$ for some generating sets $X_{\beta}\subseteq\mathfrak{g}_{\beta}$ for non-ultrashort $\beta$ and generating sets $(X_{\beta},X_{2\beta})\subseteq(\mathfrak{g}_{\beta}\mathbin{\dot{\oplus}}\mathfrak{g}_{2\beta},\mathfrak{g}_{2\beta})$ for ultrashort $\beta$. It is easy to see by considering all rank $2$ and rank $3$ irreducible root systems that $\Gamma_{\alpha}$ is closed and $\angle(\alpha,\beta)\leq\frac{\pi}{2}$ for all $\beta\in\Gamma_{\alpha}$.
@@
  • If the type of $\Phi$ is $\mathsf{BC}_{\ell}$ and $\alpha$ is ultrashort, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}^{\mathrm{us}}(G)\,U_{\alpha}$.
  • If $\alpha=\mathrm{e}_{i}+\mathrm{e}_{j}$ (using the convention $\mathrm{e}_{-i}=-\mathrm{e}_{i}$) is short and the type of $\Phi$ is $\mathsf{C}_{\ell}$ or $\mathsf{BC}_{\ell}$, including $\mathsf{B}_{2}=\mathsf{C}_{2}$, then $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,U_{2\mathrm{e}_{i}}\,U_{\mathrm{e}_{i}+\mathrm{e}_{j}}\,U_{2\mathrm{e}_{j}}$.
  • Otherwise $\operatorname{C}_{G}(Z_{\alpha})=\operatorname{C}(G)\,U_{\alpha}$.
@@
Considering saturated root subsystems of rank $2$ and $3$ containing $\alpha$ and arbitrary $\beta\in\Sigma$ (or just explicitly calculating $\Gamma_{\alpha}$ in all cases) we see that